OpenStudy (xcecex):

Help me please! A medal and fan will be awarded :) 4x/ (5-3) ---------- 2x*5+9x

OpenStudy (xcecex):

Put the answer in lowest terms.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this is not really legible use the equation tool and we can do it

OpenStudy (xcecex):

OH MY GAWD THE FIRST PERSON WHO LOOKED AT MY QUESTION

OpenStudy (xcecex):

Okay :)

OpenStudy (misty1212):

it might be that others looked but couldn't read what it was supposed to be, so didn't answer

OpenStudy (xcecex):

\[4x \div (5-3)\2x \times 5+9\]

OpenStudy (xcecex):

..? why is that little red thing right dere

OpenStudy (misty1212):

can you post a screen shot or something? this makes no sense

OpenStudy (xcecex):

okay hold on

OpenStudy (xcecex):

srry wrong link XD

OpenStudy (xcecex):

Omg forget it. Idk why it doesnt make sense it makes sense to me :l

OpenStudy (xcecex):

Its 4x divided by (5-3) Over 2x times 5 + 9x

OpenStudy (misty1212):

are those x's times signs or variables?

OpenStudy (xcecex):

Variables :l

OpenStudy (misty1212):

\[\frac{4x}{5-3}=\frac{4x}{2}=2x\]

OpenStudy (xcecex):

Okay so 2x is the numerator answer and now we have to do the denominator right?

OpenStudy (xcecex):

We dont have to answer it we just have to put it in its lowest terms

OpenStudy (misty1212):

\[2x(5x+9)=10x^2+18x\] but i am not at all sure this is what you mean

OpenStudy (xcecex):

Nvm ill just close this question thanks for trying :)